A Fraunhofer diffraction pattern is produced on a screen located 1.00m from a single slit. If a light source of wavelength 5.00 x 10^(-7) m is used and the distance from the center of the central bright fringe to the first dark fringe is 5.00 x 10^(-3) m, what is the slit width?


1
Expert's answer
2021-02-28T07:28:50-0500

Given,

The distance between the screen and the slit (D)=1.0m

Wavelength of the light source "(\lambda) = 5\times 10^{-7}m"

for the dark fringe "(y)=5.0\times 10^{-3}m"

The width of the slit "(a)=\frac{m\lambda D}{y}"

Now, substituting the values,

"a=\frac{1\times 5\times 10^{-7}\times 1}{5\times 10^{-3}}" m

"\Rightarrow a = 1.0 \times 10^{-4}" m